### v1

# AgriField3D: A Curated 3D Point Cloud and Procedural Model Dataset of Field-Grown Maize from a Diversity Panel

Abstract
The application of artificial intelligence (AI) in three-dimensional (3D) agricultural research, particularly for maize, has been limited by the scarcity of large-scale, diverse datasets. While 2D image datasets are abundant, they fail to capture essential structural details such as leaf architecture, plant volume, and spatial arrangements that 3D data provide. To address this limitation, we present AgriField3D (website), a curated dataset of 3D point clouds of field-grown maize plants from a diverse genetic panel, designed to be AI-ready for advancing agricultural research. Our dataset comprises over 1,000 high-quality point clouds collected using a Terrestrial Laser Scanner, complemented by procedural models that provide structured, parametric representations of maize plants. These procedural models, generated using Non-Uniform Rational B-Splines (NURBS) and optimized via a two-step process combining Particle Swarm Optimization (PSO) and differentiable programming, enable precise, scalable reconstructions of leaf surfaces and plant architectures. To enhance usability, we performed graph-based segmentation to isolate individual leaves and stalks, ensuring consistent labeling across all samples. We also conducted rigorous manual quality control on all datasets, correcting errors in segmentation, ensuring accurate leaf ordering, and validating metadata annotations. The dataset further includes metadata detailing plant morphology and quality, alongside multi-resolution subsampled versions (100k, 50k, 10k points) optimized for various computational needs. By integrating point cloud data of field grown plants with high-fidelity procedural models and ensuring meticulous manual validation, AgriField3D provides a comprehensive foundation for AI-driven phenotyping, plant structural analysis, and 3D applications in agricultural research.
1 Introduction
Understanding phenotypic traits is essential for improving crop yield, resilience, and sustainability in agricultural research [1, 2]. Phenotypic analysis facilitates the study of key characteristics such as plant architecture, leaf angles, and biomass allocation, which are directly linked to yield and environmental adaptability [3, 4, 5, 6, 7]. Among staple crops, maize occupies a pivotal role due to its economic importance, global cultivation, and significant contribution to food security [4, 7]. With the increasing global demand for sustainable agricultural practices, robust phenotyping methods are indispensable for advancing crop improvement programs.
The emergence of advanced 3D imaging technologies, including terrestrial LASER scanning (TLS) [8] and photogrammetry [9], has revolutionized plant phenotyping by offering detailed structural insights beyond the limitations of traditional image-based 2D methods [10, 11, 12]. 3D phenotyping enables the precise characterization of complex plant traits, such as canopy structure, volumetric growth, and organ-specific morphology. These methods have been successfully applied to diverse crops, including rubber tree forests, weeds, wheat, and rice, demonstrating their versatility in agricultural research [13, 14, 15, 16]. By capturing spatial and volumetric features with high accuracy, 3D phenotyping enables detailed plant trait analysis.
Despite these advances, 3D phenotyping poses several challenges, including data complexity, variability in plant structures, and the computational demands of processing high-resolution datasets. Machine learning (ML) and artificial intelligence (AI) have emerged as transformative tools to address these challenges, enabling automated extraction of meaningful phenotypic traits from large and complex 3D datasets [17, 18]. AI-powered methods have been employed for tasks such as predicting biomass, analyzing leaf traits, and detecting early stress responses [6]. Innovative approaches, including weakly supervised learning frameworks such as Eff-3DPSeg [18] and deep learning models for point cloud analysis [19], have further advanced organ-level segmentation and trait quantification. These technologies significantly enhance the efficiency and accuracy of 3D phenotyping but are hindered by the lack of high-quality, annotated datasets.
A key bottleneck in leveraging AI for 3D plant phenotyping is the scarcity of extensive and agronomically contextual datasets [20, 21]. Effective training and evaluation of ML models require datasets that capture the complexity of plant structures, span various growth stages, and include detailed annotations such as semantic and instance labels. Well-annotated datasets also facilitate benchmarking and development of algorithms tailored to agricultural applications, bridging the gap between theoretical advancements and real-world deployment.
In the past few years, there has been increasing interest in creating 3D plant datasets to address phenotyping needs. However, creating comprehensive real-world 3D datasets presents significant challenges, especially in agriculture. The process requires substantial resources, including field and greenhouse space, expert agricultural knowledge, and extended periods for plant growth cycles, making it both expensive and complex. To overcome these limitations, researchers have explored synthetic alternatives that can generate large-scale datasets efficiently. For example, Zarei et al. [22] utilized procedural modeling to generate 64,000 annotated sorghum plants, while Anagnostopoulou et al. [23] created 15,000 annotated mushroom images using generative models. While this virtual generation method offers a promising alternative for creating extensive datasets, the synthetic data may not fully capture the nuances of real plants, especially under field-grown conditions, potentially limiting its practical applications.
Table 1: Comparison of publicly available 3D plant datasets with annotated 3D models.
Dataset # Models (species) Modality Color Labels
Pheno4D 84 (Maize), 140 (Tomato) Laser No Semantic & Instance. Includes 49 labeled maize point clouds and 77 labeled tomato point clouds.
ROSE-X 11 (Rose) X-rays No Semantic. Voxel-level annotations with organ-level labels (stem, leaf, flower).
Single Maize Point Cloud 428 (Maize) Laser No Semantic & Instance. Includes semantic and instance annotations of stem and leaves.
Plant3D 152 (Arabidopsis), 311 (Tomato), 105 (Tobacco), 141 (Sorghum) Laser No Growth conditions. Scanned across 20–30 developmental time points under conditions like ambient light, high heat, high light, vegetative shade, and drought.
Soybean-MVS 102 (Soybean) MVS Yes Semantic. Includes 5 varieties, with semantic labels for plant organs like leaves, main stems, and stems.
PLANesT-3D 10 (Pepper), 10 (Rose), 14 (Ribes) SfM-MVS Yes Semantic & Instance. Annotated 3D color point clouds with labels for ”leaf” and ”stem,” and uniquely labeled individual leaflets.
Crops3D 83 (Tomato), 196 (Cabbage), 118 (Potato), 20 (Potato plot), 176 (Cotton), 150 (Rapeseed), 14 (Rapeseed plot), 148 (Wheat), 225 (Maize), 16 (Maize plot), 84 (Rice), SfM-MVS, Structured-Light, TLS Yes Semantic & Instance. Includes organ-level segmentation for soil, stem, leaf, tassel, fruit, pot, and more. Instance annotations available for individual plants within plots.
AgriField3D (Ours)  1,045 (Maize) TLS, Procedural models Yes Semantic & Instance. Includes 520 annotated point clouds with segmentation for leaves and stalks. Provides multiple subsampled datasets (100k, 50k, 10k) tailored for different downstream applications. High-resolution AI-ready data ideal for phenotypic analysis and ML applications.
Table 1 provides a comparison of existing publicly available 3D plant datasets, highlighting their key characteristics and limitations. These datasets vary widely in size, scanning technology, and annotation detail. The smallest dataset, ROSE-X, includes 11 rose plant models, while Crops3D is the largest with 1,230 3D plant point clouds across eight crop species. Among single-crop datasets, the Single Maize Point Cloud dataset, with 428 maize samples, represents the largest publicly available collection dedicated to a single species.
Pheno4D [24] introduces a dataset featuring 7 maize plants and 7 tomato plants. These plants were scanned daily using a 3D laser scanning system over two weeks for maize and three weeks for tomato plants, resulting in 84 point clouds for maize (with 49 labeled) and 140 point clouds for tomato (with 77 labeled). The dataset does not include color. Labels applied to the dataset include semantic and instance labels. ROSE-X [25] provides a dataset of 11 complete 3D models of real rosebush plants. These models were acquired using X-ray computed tomography (CT), offering voxel-level annotations for the entire plant shoot, with organ-level labels for stem, leaf, and flower. While the organ labels are color-coded in the point cloud representations, the dataset does not specifically include true color information. A dataset of 428 maize samples was introduced by [26], where point clouds were collected using a 3D laser scanner. These samples spanned five maize varieties (Xian Yu 335, LD145, LD502, LD586, and LD1281) and were annotated using the Label3DMaize toolkit for semantic and instance segmentation. Downsampled point clouds were processed to a uniform size of 20,480 points per sample, ensuring efficient use in training. A total of 709 plant shoot architectures from four species (Arabidopsis, Tomato, Tobacco, and Sorghum) are featured in Plant3D [27, 28, 29]. These architectures have been scanned across 20–30 developmental time points and under a variety of situations, including ambient light, high heat, high light, vegetative shade, and drought. Soybean-MVS [30] introduces a dataset consisting of 102 soybean 3D plant models reconstructed using multiple-view stereo (MVS) technology. The dataset spans the entire soybean growth period and includes five varieties. Color data was included in the point clouds (xyzRGB format). The dataset features semantic labels for plant organs, including leaves, main stems, and stems. A dataset of 34 actual plant point clouds from three species—Capsicum annuum (pepper), Rosa kordana (rose), and Ribes rubrum (ribes)—is presented by PLANesT-3D [31]. These plants were reconstructed using Structure from Motion (SfM) and MVS techniques, offering annotated 3D color point clouds. Points were labeled as either ”leaf” or ”stem,” and individual leaflets within the plants were uniquely labeled. Crops3D [32] provides a comprehensive dataset comprising 1,230 colored 3D crop point cloud samples from eight crop species including cabbage, cotton, maize, potato, rapeseed, rice, tomato, and wheat. Point clouds were acquired using a combination of TLS, Structured Light Scanning, and SfM-MVS techniques. Points were categorized into specific plant organs such as soil, stem, leaf, tassel, fruit, pot, and others, depending on the crop. The dataset includes instance annotations for individual plants within plots.
While these datasets have advanced plant-specific ML algorithms and segmentation methods, they are often limited by challenges such as low-resolution models, incomplete annotations, small dataset sizes, and restricted diversity of plant conditions. For maize, existing datasets are small in scale, lack comprehensive annotations, or fail to incorporate field-grown plants representing real-world phenotypic variability. These gaps hinder the development of robust AI-driven phenotyping solutions, particularly for traits like organ-level segmentation and trait analysis.
To overcome these limitations, we introduce AgriField3D, a curated, large-scale dataset with 1,045 high-resolution 3D point clouds of field-grown maize plants captured using TLS and additionally represented using procedural modeling techniques. This dataset was collected from the SAM diversity panel. It includes 520 annotated point clouds with semantic and instance segmentation for leaves and stalks. Additionally, AgriField3D incorporates procedurally generated maize models, which were created for each of the 520 segmented maize plants as described in Section 3.1.5. This procedural modeling approach allows for low dimensional, surface based representation of realistic maize canopies by leveraging actual plant data. The procedural models provide advantages for AI approaches including data augmentation, and synthetic-to-real transfer learning. To further support AI-driven research, AgriField3D provides multiple subsampled versions (100k, 50k, and 10k points per cloud) optimized for different computational needs.
The main contributions of AgriField3D are:
  • 1,045 quality-controlled 3D point clouds of diverse field-grown maize plants.
  • Organ-level annotations with detailed semantic and instance segmentation, enabling precise plant structure analysis.
  • AI-compatible formats, ensuring seamless integration with advanced ML models.
  • Color information, enhancing data interpretability.
  • Procedurally generated maize leaf models, providing a NURBS surface-based representation that is beneficial for downstream tasks.
  • Multiple subsampled versions (100k, 50k, and 10k points per cloud), supporting a range of computational requirements.
Our dataset is among the largest and most detailed resources for maize phenotyping, providing a robust foundation for training and validating AI-based phenotyping algorithms.
In the following sections, we detail the data processing methodologies, dataset annotation, metadata and quality control, and potential applications and highlight its significance for advancing agricultural research. By providing this dataset, we aim to catalyze the development of innovative AI-driven phenotyping solutions, paving the way for new insights into maize biology and crop improvement.
2 Dataset Description
2.1 Dataset access
The AgriField3D dataset [^1] https://baskargroup.github.io/AgriField3D/ is publicly available through the Hugging Face Datasets platform [^2] https://huggingface.co/datasets/BGLab/AgriField3D, ensuring accessibility and ease of integration into machine learning pipelines. The dataset includes high-resolution 3D point clouds and segmented plant models, organized for efficient navigation and usability. Users can download the complete dataset or select specific components tailored to their research needs. Key features include:
  • Point Clouds: Point cloud data of over 1,000 maize plants at multiple subsampled resolutions ($10,000$, $50,000$, and $100,000$ points) to accommodate various computational requirements.
  • Segmentation: Detailed segmented models of 520 plants, including individual leaves and stalks, with subsampled resolutions ($100,000$, $50,000$, and $10,000$ points). Each leaf is consistently color-coded based on its vertical position, and stalks are assigned distinct colors, facilitating organ-specific visualization and analysis.
  • Metadata: A comprehensive metadata file (Metadata.xlsx) provides detailed annotations for each plant, ensuring clarity and contextual understanding of the dataset.
  • Documentation: A detailed user guide is included in the repository, offering step-by-step instructions for integrating the dataset with tools such as PyTorch, TensorFlow, and Scikit-learn, enabling streamlined machine learning applications.
To enhance the dataset’s usability, the following reconstructed outputs of integration of the segmented dataset with the procedural model explained in Section 3.1.5 are also included:
  • STL Files: These files include the 3D reconstructed surfaces, compatible with visualization and simulation tools.
  • DAT Files: These files provide the control point information for the NURBS surfaces, enabling further refinements and integration into CAD workflows.
Figure 1: Nine examples of maize plant point clouds from the dataset, showcasing the original data. These images highlight the diverse morphologies of maize plants captured using terrestrial laser scanning.
The Hugging Face repository supports both full dataset downloads and selective access to subsets, such as segmented plants or point clouds at specific resolutions. This flexibility allows users to tailor the dataset to their specific requirements. For any inquiries or assistance in accessing specific components, users are encouraged to contact the corresponding authors via the email addresses provided in this publication.
2.2 Dataset Composition
The dataset comprises over 1000 high-resolution point clouds, each capturing detailed 3D information of individual maize plants. The dataset represents a diverse range of genetic backgrounds and provides the following key components:
  • High-resolution point clouds for various maize plants.
  • Around 520 segmented plants into stalks and leaves with specific color coding, ordered from bottom to top leaves.
  • Metadata detailing plant quality, leaf count, presence of tassels and ears, and other critical attributes.
  • Multiple resolutions to support diverse AI applications.
Figure 1 showcases nine example maize plants from the dataset, highlighting the original data without segmentation and the diversity in plant morphology.
2.3 Data Collection
Field-grown maize plants harvested at about the time of anthesis were selected for data collection from the maize SAM (shoot apical meristem) diversity panel [33] at Iowa State University’s Curtis Farm. These lines were planted in 4-row plots at the Iowa State University Woodruff Farm in Ames, Iowa, in the year 2021. Each plot measured 3.04 meters in row length, with plants spaced approximately 15.24 cm apart within rows and 76.2 cm apart between rows, resulting in a planting density of around 84,000 plants per hectare. All plants were collected from the field at the time of anthesis and immediately brought to indoor environment for imaging and 3D scanning. Each plant was positioned upright using specialized holders to maintain a natural posture, with barcodes scanned to link each plant to its variety tag. Corresponding tag names were printed on white sheets and placed adjacent to the plants for precise labeling and consistent organization.
The 3D data were captured using a Faro Focus S350 Scanner, depicted in Figure 2(a), which is equipped with an angular resolution of 0.011 degrees, enabling precise point spacing of approximately 1.5 mm at a 10-meter range. The scanner has the capacity to acquire up to 700 million points at a rate of 1 million points per second, ensuring high-resolution outputs suitable for detailed phenotypic analysis. Plants were arranged in batches of eight around the scanner in a circular configuration, as shown in Figure 3, allowing simultaneous capture of multiple samples. This setup minimized overlaps between plants and ensured most surfaces were within the scanner’s line of sight. The scanner height and settings were dynamically adjusted based on plant height and morphology, ensuring that critical structural features, including upper leaves, stalks, tassels, and ears, were precisely captured. Figure 2(b) provides a visual representation of the scanning setup, showcasing the Faro scanner in operation while capturing maize plants arranged in a circular configuration. The controlled indoor environment eliminated external variables such as wind and inconsistent lighting, ensuring uniformity and accuracy across the dataset. Scans were conducted batch by batch, until the full dataset was acquired.
Figure 2: (a) The Faro Focus S350 Scanner is used for collecting 3D data of maize plants. This advanced scanner ensures precise and high-resolution point cloud generation, which is critical for phenotypic analysis. (b) Visualization of the scanning process using the Faro scanner. Maize plants are arranged in a circular configuration, allowing simultaneous high-resolution capture of multiple samples in a controlled indoor environment.
Figure 3: Point cloud visualization of maize plants. (a) The circular arrangement of eight maize plants mounted on a specialized platform, each labeled with unique identifiers. (b) Detailed view of a single maize plant showing characteristic leaf architecture and stem structure.
2.4 Data Processing
The raw 3D scan data were processed into detailed 3D point clouds using FARO SCENE software, which integrated RGB color values to the outputs. The area of interest, containing only the plants, was then cropped from the scans to exclude extraneous elements. To reduce noise, statistical outlier removal was applied using both global and local point-to-point distance distributions, effectively eliminating artifacts caused by environmental interference or scanner irregularities. This preprocessing ensured high-quality point clouds with precise structural and color details, suitable for downstream phenotypic analyses. We detail these steps further below.
To separate individual plants, CloudCompare software was used. This process referred to as plant isolation, involved using the crop tool in CloudCompare to extract each plant from the circular arrangement. Identification and naming of the individual plants were facilitated by the paper labels placed in front of each plant during scanning. These labels ensured accurate association between physical plants and their digital representations. Once isolated, each plant was saved as an individual ‘.ply‘ file with proper naming conventions. This plant isolation process preserved the structural integrity of the plants while ensuring readiness for downstream analysis.
Following this, a quality control process was implemented. Extraneous elements such as walls, tables, and personnel inadvertently captured during scanning were removed. Each set of maize point clouds was manually inspected to verify its quality, ensuring that plant structures were intact, free from artifacts, and properly cropped. Multiple rounds of review were conducted to maintain accuracy and reliability across the dataset. Additionally, a metadata file was developed to document quality assessments and plant-specific characteristics, including the presence of tassels, maize ears, leaf count, and resolution. These metadata files provide a useful context for researchers, facilitating targeted analyses and ensuring the dataset’s suitability for diverse applications, including AI model training, phenotypic studies, and agricultural research.
3 Data Annotation
3.1 Data Segmentation
The segmentation of maize plant point cloud data is a key contribution for enabling phenotypic analysis and machine learning applications. Automated graph-based segmentation methods were combined with manual checks in this procedure. We report precise separation of plant components, including leaves and stalks, while maintaining a high standard of quality.
3.1.1 Segmentation Workflow
The initial segmentation was performed using a graph-based approach. This automated process included the following steps:
First, the structural skeleton of the point cloud was extracted through a process known as skeletonization. This step utilized Laplacian-based contraction with adjusted parameters to generalize across different maize examples, effectively identifying the primary stalk and branching patterns. Following this, stalk node identification was performed to detect the top and bottom nodes of the maize stalk based on z-coordinate extremes, with a center bias parameter applied to prioritize median x and y coordinates.
Next, the shortest path in the maize skeleton between the identified top and bottom nodes was calculated. Nodes along this path were classified as stalk nodes, while other nodes were classified as leaf nodes. To enhance segmentation accuracy, stalk nodes were temporarily removed from the skeleton graph.
The segmentation of leaves was then performed by identifying each connected component remaining in the skeleton graph after the removal of stalk nodes. Connected components with a small number of nodes were excluded to avoid noise. Subsequently, each point in the maize point cloud was assigned to its corresponding leaf or stalk based on proximity to these connected components.
Finally, points were colored with RGB values corresponding to their associated leaf or stalk, generating a point cloud file that visually represents the segmented maize plant.
3.1.2 Manual Inspection and Quality Control
Following the graph based process, we performed manual inspection and editing of the post-processed data. This manual refinement process was conducted using a Python-based preprocessing script in combination with the CloudCompare (CC) application.
Initially, each leaf and the stalk were extracted into separate .ply files for individual inspection and correction. This separation allowed for focused evaluation and adjustment of specific plant components. During this phase, common segmentation errors were addressed. For instance, misclassified points, such as parts of a leaf erroneously grouped with the stalk, were carefully segmented out and merged with the correct component. Similarly, interleaf overlaps, where points from one leaf were mistakenly grouped with a neighboring leaf, were rectified by accurately reassigning those points. To maintain consistency, the leaf numbering was reviewed and reordered from the bottom to the top of the plant. This step ensured a clear structure in the dataset.
3.1.3 Recoloring and Merging
Following segmentation, a recoloring and merging process was implemented. Each plant component was recolored to adhere to a consistent color scheme. For example, the first leaf in every plant was assigned the same color, followed by the second leaf, and so on, with the stalk uniformly colored across all samples. This uniformity enhanced the dataset’s usability and interoperability. Table 2 details the color ordering of the leaves and stalks.
After recoloring, the corrected components were merged into a single .ply file for each plant. This consolidated format preserved the color labels, allowing for easy identification of individual leaves and the stalk within each plant. The merged files provided a cohesive representation of the plant structure while retaining the detailed segmentation information.
Table 2: Color Coding for Plant Components
Component Color Name RGB Values Hex Code
Leaf 1 Bright Blue [0, 0, 255] #$0$FF
Leaf 2 Bright Red [255, 0, 0] #FF0000
Leaf 3 Indigo [75, 0, 130] #4B0082
Leaf 4 Deep Pink [255, 20, 147] #FF1493
Leaf 5 Dark Goldenrod [184, 134, 11] #B8860B
Leaf 6 Purple [128, 0, 128] #800080
Leaf 7 Orange Red [255, 69, 0] #FF4500
Leaf 8 Dark Blue [0, 0, 139] #00008B
Leaf 9 Saddle Brown [139, 69, 19] #8B4513
Leaf 10 Bright Green [0, 255, 0] #00FF00
Leaf 11 Dark Orange [255, 140, 0] #FF8C00
Leaf 12 Sea Green [46, 139, 87] #2E8B57
Leaf 13 Gold [255, 215, 0] #FFD700
Leaf 14 Dark Slate Blue [72, 61, 139] #483D8B
Leaf 15 Steel Blue [70, 130, 180] #4682B4
Leaf 16 Brown [165, 42, 42] #A52A2A
Leaf 17 Teal [0, 128, 128] #008080
Leaf 18 Bright Orange [255, 165, 0] #FFA500
Stalk Black [0, 0, 0] #000000
3.1.4 Segmented Dataset
The segmentation and refinement processes were applied to 520 plants.
Figure 4: Nine examples of segmented maize plant point clouds from the dataset, showcasing the color-coded segmentation data.
Figure 4 showcases nine examples of the segmented maize plant point clouds, highlighting the color-coded segmentation data among the 520 segmented plants in the dataset.
3.1.5 Procedural Model
Procedural modeling is a computational approach to generating structured 3D representations from raw data by applying algorithmic rules and mathematical functions. Unlike raw point clouds, which are unstructured and can be noisy or incomplete, procedural models create parametric surfaces that provide a compact, scalable, and editable representation of plant structures. This method enables precise reconstructions that preserve geometric accuracy while allowing for controlled variations, making it particularly useful for synthetic data augmentation, structural analysis, and AI-driven phenotyping applications. Procedural modeling is especially valuable for plant phenotyping, where capturing complex leaf curvature and growth patterns is essential for trait analysis and computational simulations. The segmented dataset of 520 maize plant point clouds was used as input for a procedural modeling approach developed in Hadadi et al. [34]. This method enabled the high-fidelity reconstruction of 3D surfaces that accurately capture the shape and structure of individual maize leaves. By leveraging this procedural model, each segmented plant was processed to NURBS-based surfaces. The procedural model ensured that the reconstructed surfaces were closely aligned with the original segmented point clouds. Figure 5 and Figure 6 illustrate examples of these reconstructions, showcasing the transition from raw point cloud data to the final NURBS surface representation.
Figure 5: Processing of B73 maize plant. (a) Raw data, (b) segmented data, and (c) final NURBS surface.
Figure 6: Processing of Mo17 maize plant. (a) Raw data, (b) segmented data, and (c) final NURBS surface.
To enhance the usability of this dataset, the reconstructed outputs are made publicly available in the following formats:
  • STL Files: Containing 3D reconstructed surfaces, suitable for visualization and simulation applications.
  • DAT Files: Providing control point data for the NURBS surfaces, allowing further refinements and integration into CAD workflows.
These additional files complement the segmented point clouds, bridging raw data with high-fidelity surface models that can be utilized for diverse research applications.
Figure 7: NURBS surface reconstruction using the procedural model for four diverse maize plants.
Figure 7 illustrates the NURBS surface reconstructions derived from the procedural modeling approach applied to four distinct maize plants. The examples highlight the diversity in maize leaf structures across maize plant varieties. The process accurately captures the complex geometries of maize leaves, including their curvature and edges.
4 Data Downsampling
Upon analyzing the dataset, substantial variability was observed in the number of points per point cloud. The point counts ranged from as low as 150K points to as high as 2.6 million points, with most point clouds centered around 600K points. This variability, while reflective of the diversity in plant structures, may pose challenges for downstream tasks requiring uniform point counts.
Figure 8 provides an overview of the original point cloud dataset before any downsampling or modifications. The histograms highlight two key aspects of the dataset:
  • Point Count Distribution (Left Plot): The first histogram depicts the distribution of the number of points per plant, scaled to millions. The dataset exhibits a wide range of point counts across the plants, reflecting the diversity in plant structure and size. Most plants have a point count within a specific range, with fewer plants having extremely high or low counts.
  • Point to Point Distance (Right Plot): The second histogram shows the distribution of the mean nearest neighbor distances, converted to millimeters. This measure captures the spatial density of points within each plant’s point cloud. A narrower spread in this distribution indicates relatively consistent point density across the dataset.
The left panel demonstrates the scale and resolution of the original dataset, while the right panel highlights the structural fidelity retained in the point cloud representation. Such metrics are crucial for understanding how well the dataset represents real-world plant morphology and for determining appropriate preprocessing steps for downstream applications.
Figure 8: Histograms summarizing the original point cloud dataset before downsampling. The left histogram illustrates the distribution of the number of points per plant, scaled to millions, highlighting the variability in point cloud sizes across the dataset. The right histogram shows the distribution of mean nearest neighbor distances, converted to millimeters, capturing the spatial density of points within each plant’s point cloud. Together, these histograms provide insights into the scale and resolution of the dataset, emphasizing its diversity and structural fidelity.
To address this variability, downsampled versions of the dataset were created with fixed point counts $N$. These standardized datasets ensure compatibility with computational frameworks and machine-learning pipelines that require consistent input sizes. Additionally, downsampling reduces computational overhead, broadening the dataset’s usability for researchers with varying computational resources. Three target point counts were chosen to meet diverse research needs, including $100,000$, $50,000$, and $10,000$ points. These resolutions accommodate applications ranging from resource-intensive tasks requiring high-resolution data to computationally efficient analyses.
Several downsampling techniques were evaluated to identify the most suitable method for downsampling. We used Open3D[35] for point cloud processing including voxel grid downsampling, Poisson disk sampling, farthest point sampling, and random downsampling. In our implementation, we combine two downsampling steps. First, a voxel grid downsampling procedure is applied, which uniformly partitions the space and reduces the point count while preserving the overall structure of the plants. The algorithm iteratively increases the voxel size until the downsampled point cloud has just more than the target $N$ points. If the resulting point cloud still exceeds $N$, a fine-tuning step is performed using random sampling to exactly reach the desired count. This combined approach balances computational efficiency with data quality and ensures precise point counts. Additionally, to further enhance usability, we have provided a Python script to allow researchers to generate custom downsampled datasets with any desired point count $N$.
Figure 9: Comparison of histograms showing the mean point to point distances for the dataset downsampled to 100K points, 50K points and 10K points.
Figure 9 presents the histograms of mean nearest neighbor distances for the three downsampled versions. As the point count decreases, the mean point-to-point distances increase, reflecting the reduced resolution. These visualizations highlight the trade-off between dataset size and structural detail, enabling researchers to choose the resolution best suited to their computational and analytical needs.
4.1 Metadata
The metadata file contains the following columns:
  • TagName: The unique identifier for the plant, associated with its genotype through a separate look-up table.
  • IsConnected: Indicates whether the plant’s stem and leaves are fully connected with no gaps (Yes/No).
  • IsEasyToCountLeaves: Categorizes the ease of counting leaves as easy, moderate, or hard.
  • AreLeavesTouching: Specifies whether the leaves of the plant are touching each other (Yes/No).
  • NumberOfLeaves: Records the total number of leaves present on the plant.
  • HasColor: Indicates the presence of RGB color data in the point cloud (Yes/No).
  • HasEar: Indicates whether the plant has an ear (Yes/No).
  • HasTassel: Indicates whether the plant has a tassel (Yes/No).
  • Size: Specifies the number of points in the point cloud.
  • Comment: Includes any additional notes or observations relevant to the plant or the quality of its point cloud.
Column Name Value
TagName 0433
IsConnected Yes
IsEasyToCountLeaves Easy
AreLeavesTouching No
NumberOfLeaves 15
HasColor Yes
HasEar Yes
HasTassel Yes
Size 1,558,177 points
Comment This is a good plant.
Figure 10: Side-by-side representation of a maize plant point cloud and its corresponding metadata table. The point cloud is visualized on the left, while the table on the right provides detailed metadata annotations, enabling researchers to interpret the dataset effectively for downstream analyses.
The metadata not only provides a detailed description of each plant’s structural and phenotypic characteristics but also allows for targeted data selection. For example, researchers can quickly identify plants with a specific number of leaves or those with complete color information to suit their experimental requirements.
5 Potential Applications
The curated 3D point cloud dataset of maize plants offers numerous opportunities for advancing agricultural research and crop improvement. We identify some promising applications below.
Feature Extraction and Phenotypic Analysis:
The dataset’s detailed segmentation and annotations support the extraction of critical phenotypic traits, including the number of leaves, leaf length, width, and curvature, as well as stalk dimensions. These features can be used to quantify plant architecture, model growth patterns, and analyze genotype-phenotype relationships [36]. The accurate segmentation of leaves and stalks further facilitates studies on plant competition, light interception, and resource allocation within the canopy [37].
Shape Representation and 3D Modeling:
The high-resolution 3D point clouds serve as valuable benchmarks for developing and validating advanced neural shape representation models, including implicit neural fields and sparse representations [38]. These models can capture the structural complexity of plants with reduced computational overhead, making them valuable for large-scale phenotyping studies. Researchers can use the dataset to train and evaluate machine learning models that generalize across diverse genotypes and growth conditions, improving the robustness of predictive algorithms [39].
Virtual Phenotyping and Computational Plant Modeling:
The dataset supports the exploration of novel methods for 3D shape analysis, including applications in virtual phenotyping and computational plant modeling. For example, the integration of neural representations with segmented point clouds can enable the simulation of plant responses to environmental stimuli, such as light and wind [40]. These simulations can contribute to optimizing crop designs for improved yield and resource efficiency.
Agricultural Technology Development:
In addition to scientific research, the dataset has potential applications in agricultural technology development. It can be used to train and validate algorithms for autonomous systems, such as robotic harvesting and precision agriculture tools [41]. These systems rely on accurate 3D representations of plants for navigation, crop monitoring, and targeted interventions.
6 Conclusions
This study presents a high-resolution 3D point cloud dataset of maize plants. By capturing detailed structural features and providing comprehensive metadata, this dataset bridges a critical gap in agricultural research, enabling precise and scalable analyses of plant architecture, morphology, and genotype-phenotype relationships. The dataset’s segmentation, annotation, and standardized downsampling cater to a wide range of downstream applications, from feature extraction and shape representation to the development of machine learning models for plant phenotyping. Its versatility makes it a valuable tool not only for phenomics but also for agricultural technology development, such as precision farming and autonomous systems.
While the dataset represents a significant step forward in providing high-resolution 3D point clouds for plant phenomics, limitations must be acknowledged. A key limitation is the focus on a single growth stage. The dataset captures plants at vegetative maturity, providing valuable insights into their structural characteristics at this stage. However, it does not encompass the full developmental cycle of maize, such as seedling, flowering, or senescence stages. This restricts the scope of analyses to specific phenotypic traits relevant to vegetative maturity.
Future work will aim to address these limitations by expanding the dataset and incorporating additional metadata fields to enhance its utility. Specifically, future versions of the dataset will include data from multiple growth stages, enabling researchers to study dynamic phenotypic changes throughout the plant lifecycle. These additions will provide a more comprehensive resource for investigating developmental processes and genotype-phenotype interactions over time.
By providing this dataset openly to the research community, we aim to accelerate progress in plant science and crop improvement studies. The integration of detailed metadata further enhances its usability, allowing researchers to tailor their analyses to specific traits and research objectives. We encourage researchers to explore this dataset in diverse agricultural contexts, from studying plant responses to environmental conditions to improving crop yield and sustainability. Future expansions of this resource, including additional growth stages and environmental data, will further amplify its impact, fostering innovation and collaboration across the field.
Acknowledgments
This work is supported by the AI Research Institutes program [AI Institute for Resilient Agriculture (AIIRA), Award No. 2021-67021-35329] from the National Science Foundation and U.S. Department of Agriculture’s National Institute of Food and Agriculture. We acknowledge support from Iowa State University Plant Science Institute. YL was supported in part by NSF grants (DBI-1661475 and IOS-1842097) to PSS and others and a scholarship from the China Scholarship Council.